OpenStudy (anonymous):

Evaluate the polynomial two ways:by substituting in the given value of x, and by using synthetic division: Find P(-4) for P(x)=x^4+2x^3-13x^2-39

OpenStudy (dumbcow):

P(-4) = (-4)^4 +2*(-4)^3 -13*(-4)^2 -39 =256-128-208-39 =-119 synthetic division: -4 | 1 2 -13 0 -39 -4 8 20 -80 ------------------------- 1 -2 -5 20 -119

OpenStudy (anonymous):

Thanks so much.

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

Do you have any idea of how to work this one? approximate to two decimal places, find the x-intercepts, local maximum and local minimum: p(x)=x^3+2x^2-7x-5

OpenStudy (dumbcow):

it doesnt factor so you need to graph the function or plot points to identify those points